Circular No. 6516

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J2058+42 A. J. Castro-Tirado, Laboratorio de Astrofísica Espacial y Fisica Fun-

damental, Madrid, communicates: “We have examined the 8′ RXTE error box of the x-ray pulsar GRO J2058+42 (IAUC 6514) on a deep Schmidt plate taken by K. Birkle at the German-Spanish Calar Alto Observatory on 1995 Oct. 15 in response to the report of the outburst on IAUC 6238. On that date, the pulsed signal was about 70 mCrab (C. A. Wilson, private communication.). A new object of mag ∼ 18.5 not present on the corresponding Palomar Sky Survey plate is seen at α = 20^{h}58^{m}58^{s}.0, δ = +41^{o}43′06′′ (equinox 2000.0, uncertainty ±1′′). A second object of mag ∼ 19.5 that seems to be variable by ∼ 0.5 mag is also seen at α = 20^{h}59^{m}07^{s}.7, δ = +41^{o}44′46′′. A finding chart can be obtained via anonymous ftp at

laeff.esa.es (under /pub/users/ajct).”

SUPERNOVA 1996bw IN NGC 664 A. V. Filippenko, D. C. Leonard and W. Ho, University of California

at Berkeley, report that inspection of a CCD spectrum (range 320–980 nm, resolution 0.7 nm) obtained on Dec. 7 with the 3-m Shane reflector at the Lick Observatory reveals that SN 1996bw is a young type II (rather than type Ib; cf. IAUC 6512) supernova. The nearly featureless continuum is very blue, but weak, broad absorption lines of Hβ and Hγ are clearly visible. There may be a low-contrast, broad Hα emission line; the corresponding absorption component is very weak. M. Cavagna and F. Manca, Sormano, Italy, have obtained the position α = 1^{h}43^{m}44^{s}.58, δ = +4^{o}13′19′′.7 (equinox 2000.0). G. J. Garradd, Loomberah, N.S.W., reports end figures 44^{s}.68, 20′′.0. D. di Cicco, Sudbury, MA, reports end figures 44^{s}.48, 19′′.8 (with unfiltered CCD mag = 17.2–17.5 on Dec. 3.2 UT).

SUPERNOVA 1956C = 1988 XB_{5} G. V. Williams, Center for Astrophysics, reports: “SN 1956C appears

to be the minor planet 1988 XB_{5} (5 oppositions, 1981–1996) near its stationary point. The object was discovered on a plate taken 1956 Feb. 10 for the Palomar Sky Survey that was subsequently rejected (Kowal et al. 1971, PASP 83, 307). The position of the galaxy is given as α = 12^{h}43^{m}.1, δ = +3^{o}49′ (equinox 2000.0), the presumed supernova being of m_{pg} = 17.6 and offset by 5′′ east, 27′′ south. The predicted positions of 1988 XB_{5} for 1956 Feb. 10.0 and 11.0 UT are α = 12^{h}43^{m}.18, δ = +3^{o}47′.4 and 12^{h}43^{m}.26, +3^{o}49′.5, respectively, the predicted magnitude being V = 17.0.”

1996 December 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden